OpenStudy (anonymous):

For the following system of equations, what is the x-value of the solution? –x + 2y = 6 6y = x + 18

OpenStudy (anonymous):

8 for the first one

OpenStudy (anonymous):

thanks

OpenStudy (abb0t):

divide the second equation by 6 to get your Y. substitute and solve for x. Which someone has apparently already found for you. But you might want to do the work just to see how it was done and whether their answer matches yours. Cheers.

OpenStudy (anonymous):

but there's something in the directions of substituting the y in the second equation for the y= in the first one

OpenStudy (abb0t):

\[y = \frac{ x }{ 6 }+\frac{ 18 }{ 6 } = \frac{ x }{ 6 } + 3 \]

OpenStudy (anonymous):

how'd you get that?

OpenStudy (abb0t):

I divided the 2nd equation by 6 to get y by itself so that I could use that "y" function to substitute to the top. That way, you just plug in your 2nd equation (which is composed of ONLY x's) to solve for x. Remember that you can combine like terms.

OpenStudy (anonymous):

so i then subtract the 3 and then multiply both sides by 6?

OpenStudy (anonymous):

and get -18 as my answer

OpenStudy (abb0t):

x should be: \[\frac{ 9 }{ 2 }\]

OpenStudy (abb0t):

\[-x + 2(\frac{ x }{ 6 } + 3) = 6\]